A new Neighborhood Activity ComplexNeighborhood Activity Complex is being built in Del RioDel Rio. The perimeter of the rectangular playing field is 458 yards. The length of the field is 3yards less than tripletriple the width. What are the dimensions of the playing​ field?
Answer by ikleyn(52781)

.
Let W be the width, in yards.  Then the length is L = 3W - 3 yards.


For the perimeter, you have this equation


2W + 2*(3W - 3) = 458.


Simplify and solve for W :


2W + 6W - 6 = 458,

8W = 458 + 6

8W = 464

W = 464/8 = 58.


The width is 58 yards.  The length is  3*58 - 3 = 171 yards.


CHECK.  2*58 + 2*171 = 458 yards.  ! Correct !

Solved.

-----------------

Hello,  please pay attention that your input system produces errors:  it doubles some names.

I am not very happy to decipher it.  I think that other tutors are not very happy,  too.

Fix it or inform your managers.


Also notice that the first phrase of this post can be omitted without any loss of meaning for the entire problem itself.


It is typical bla-bla-bla,  which is a routine for 80% of problem formulations in USA school Math.
Without it,  the problem would be  MUCH  BETTER.